# PISA-Bench: The PISA Index as a Multilingual and Multimodal Metric for the Evaluation of Vision-Language Models

### v1

@@ 1.  Introduction
Large language models have recently made remarkable progress, demonstrating human-like abilities in tasks such as commonsense question answering (Sun et al., 2024; Chen et al., 2024; Toroghi et al., 2024) or mathematical reasoning (Wang et al., 2025; Parashar et al., 2025; Dong et al., 2025). Despite these advances, significant performance differences remain across languages, even in language-agnostic domains such as mathematics.(Wu et al., 2025; Xu et al., 2025). Further, there are similar disparities across modalities, e.g., where models perform better on visual reasoning tasks when textual information such as image captions or OCR-extracted content is used alongside the image (Lu et al., 2024a; Yue et al., 2024a). Developing models capable of reasoning over both images and text in multiple languages is essential to mitigate the current dominance of English-centric systems (Zhu et al., 2024). Such models should be able to, for instance, identify and combine visual and textual information in any language to solve complex reasoning tasks.
@@
To address these issues, we introduce PISA-Bench, a benchmark derived from examples of the PISA tests, an international assessment of student competencies. The PISA test is a large-scale international study conducted by the OECD that evaluates the knowledge and skills of 15-year-old students in reading, mathematics, and science to assess how effectively education systems prepare them for real-world challenges (OECD, 2025).
@@ 2.1.  Stage 1: Collection of PISA Tests
We show the extracted types exemplarily in Figure˜1 (cf. 2 - Modular Information Extraction). We extract instruction, image, question, and answer options, which is, in this case, a free-form answer generation. At last, we will also extract the solution section of the original test for the gold answer. To ensure consistency and self-containment, we use GPT-4o (OpenAI et al., 2024) to (1) generate possibly missing multiple-choice options and (2) rephrase questions so that they can be answered either by selecting a multiple-choice option or by generating a free-form response. We show this prompt used for this step in in the  Appendix˜A. This allows us to evaluate our benchmark in three settings: (1) log-likelihood-based ranking of answer options, as commonly implemented in the LM Evaluation Harness (Gao et al., 2024); and free-form answer generation evaluated using either (2) string matching or (3) more advanced techniques such as LLM-as-a-judge (Zheng et al., 2023) or sentence similarity (Zhang et al., 2020).
@@ 3.  Translation Validation
As described in Section˜2.4, we translated the English source material into five languages (German, Spanish, French, Italian, and Chinese). We validate the translation quality of PISA-Bench using automatic evaluation and human verification to ensure a reliable multilingual benchmark. For automatic validation, we use two metrics: the WMT23 COMET-KIWI Rei et al. (2022) and the GEMBA-MQM Kocmi and Federmann (2023) using GPT-4. For human verification, we work with a native speaker in the respective language who verifies 50 random examples for linguistic accuracy. This quality assurance process aims to ensure that all translations in PISA-Bench are both linguistically accurate and semantically faithful to the original content.
@@ 4.  Experimental Setup
We consider the following models for our experiments: Qwen2.5-VL (3B – 72B) (Bai et al., 2025), Qwen3-VL (Team, 2025), Gemma-3 (4B – 27B) (Team et al., 2025), LLaVA (7B – 34B) (Liu et al., 2023), and Idefics3 (8B) (Laurençon et al., 2024) to cover different architectural concepts, pretraining objectives and model sizes. We further include GPT-4o and Claude-3.5-Haiku (Anthropic, 2025) as proprietary models.
@@ 5.  Evaluation Results
In Table˜4, we compare PISA-Bench with commonly known benchmarks MMLU (Hendrycks et al., 2021) and MMMU (Yue et al., 2024b). We observe that our benchmark is significantly more difficult compared to MMLU for the models investigated, e.g., we find a performance of 68.5 on MMLU using Qwen2.5-VL (7B), whereas the corresponding performance on PISA-Bench is only 52.4. However, we note that MMLU is a text-only benchmark, emphasizing that modalities other than text is more difficult for current language models.
@@ 6.  Related Work
Recent progress in large vision-language models (LVLMs) has been closely tied to the development of evaluation benchmarks. Early benchmarks primarily assessed visual perception and image understanding Fu et al. (2023); Liu et al. (2024a, 2023); Meng et al. (2024), often restricting evaluation to multiple-choice or short-form VQA tasks. More recently, these benchmarks have been extended to more general areas such as cognition and reasoning, for example, benchmarks such as M3Exam (Zhang et al., 2023), SOK-Bench (Wang et al., 2024a), MathVista (Lu et al., 2024b), or VL-ICL-Bench (Zong et al., 2025). Further, benchmarks like MMDU Liu et al. (2024b), adopt open-ended questions combined with LLM-as-a-judge evaluation, while MMMU-pro Yue et al. (2024b) propose unifying text and images into a single visual representation.
@@
In the domain of multilingual benchmarks, authors often begin with English datasets and translate them into other languages, such as XNLI (Conneau et al., 2018) or XCOPA (Ponti et al., 2020). More recent works utilize multilingual language models to translate the original test set, such as HumanEval-XL (Peng et al., 2024) or mHumanEval (Raihan et al., 2025). However, while this approach provides broad coverage, it may propagate cultural or linguistic biases through the translation using language models Shi et al. (2022). P-MMEval (Zhang et al., 2024) and BenchMAX (Huang et al., 2025) address this issue by using parallel corpora to fairly assess cross-lingual capabilities, disentangling cultural knowledge from a language model’s translation ability.
@@
Many multilingual LVLM benchmarks evaluate the general natural language and image understanding capabilities of vision-language models such as xGQA (Pfeiffer et al., 2021), GEM (Su et al., 2021), and MaXM (Changpinyo et al., 2022). More recent benchmarks extend this line of work to reasoning and cognition, including M3Exam (Zhang et al., 2023), EXAM-V (Das et al., 2024), or M5-VGR (Schneider and Sitaram, 2024). Others focus on culture-specific reasoning, such as ALM-bench (Vayani et al., 2024) and CVQA (Romero et al., 2024), or domain-specific reasoning, such as medical reasoning in WorldMedQA-V(Matos et al., 2024). To mitigate the challenges of translating datasets into target languages, parallel benchmarks such as M4U (Wang et al., 2024b), PM4Bench (Gao et al., 2025), and XT-VQA (Yu et al., 2024) have been proposed.

### v2

@@ 1.  Introduction
Large language models have recently made remarkable progress, demonstrating human-like abilities in tasks such as commonsense question answering (sun2024surveyreasoningfoundationmodels; chen2024llmbasedmultihopquestionanswering; toroghi-etal-2024-right) or mathematical reasoning (wang2025surveylargelanguagemodels; parashar2025inferencetimecomputationsllmreasoning; dong2025scalablellmmathreasoning). Despite these advances, significant performance differences remain across languages, even in language-agnostic domains such as mathematics.(wu2025bitterlessonlearned2000; Xu_2025). Further, there are similar disparities across modalities, e.g., where models perform better on visual reasoning tasks when textual information such as image captions or OCR-extracted content is used alongside the image (lu2024mathvista; yue2024mmmumassivemultidisciplinemultimodal). Developing models capable of reasoning over both images and text in multiple languages is essential to mitigate the current dominance of English-centric systems (zhu2024multilinguallargelanguagemodels). Such models should be able to, for instance, identify and combine visual and textual information in any language to solve complex reasoning tasks.
@@
To address these issues, we introduce PISA-Bench, a benchmark derived from examples of the PISA tests, an international assessment of student competencies. The PISA test is a large-scale international study conducted by the OECD that evaluates the knowledge and skills of 15-year-old students in reading, mathematics, and science to assess how effectively education systems prepare them for real-world challenges (oecd_pisa).
@@ 2.1.  Stage 1: Collection of PISA Tests
We show the extracted types exemplarily in Figure˜1 (cf. 2 - Modular Information Extraction). We extract instruction, image, question, and answer options, which is, in this case, a free-form answer generation. At last, we will also extract the solution section of the original test for the gold answer. To ensure consistency and self-containment, we use GPT-4o (openai2024gpt4ocard) to (1) generate possibly missing multiple-choice options and (2) rephrase questions so that they can be answered either by selecting a multiple-choice option or by generating a free-form response. We show this prompt used for this step in in the  Appendix˜A. This allows us to evaluate our benchmark in three settings: (1) log-likelihood-based ranking of answer options, as commonly implemented in the LM Evaluation Harness (eval-harness); and free-form answer generation evaluated using either (2) string matching or (3) more advanced techniques such as LLM-as-a-judge (zheng2023judgingllmasajudgemtbenchchatbot) or sentence similarity (zhang2020bertscoreevaluatingtextgeneration).
@@ 3.  Translation Validation
As described in Section˜2.4, we translated the English source material into five languages (German, Spanish, French, Italian, and Chinese). We validate the translation quality of PISA-Bench using automatic evaluation and human verification to ensure a reliable multilingual benchmark. For automatic validation, we use two metrics: the WMT23 COMET-KIWI rei2022cometkiwiistunbabel2022submission and the GEMBA-MQM kocmi2023gembamqmdetectingtranslationquality using GPT-4. For human verification, we work with a native speaker in the respective language who verifies 50 random examples for linguistic accuracy. This quality assurance process aims to ensure that all translations in PISA-Bench are both linguistically accurate and semantically faithful to the original content.
@@ 4.  Experimental Setup
We consider the following models for our experiments: Qwen2.5-VL (3B – 72B) (bai2025qwen25vltechnicalreport), Qwen3-VL (qwen3technicalreport), Gemma-3 (4B – 27B) (gemmateam2025gemma3technicalreport), LLaVA (7B – 34B) (liu2023visualinstructiontuning), and Idefics3 (8B) (laurençon2024mattersbuildingvisionlanguagemodels) to cover different architectural concepts, pretraining objectives and model sizes. We further include GPT-4o and Claude-3.5-Haiku (anthropic2025_claude35haiku) as proprietary models.
@@ 5.  Evaluation Results
SmolLM Family
SmolVLM2-256M-Video-Instruct 33.9±4.3 36.9±4.4 41.8±4.5 41.8±4.5 44.3±4.5 31.1±4.2 38.3 +5.3
SmolVLM2-500M-Video-Instruct 29.8±4.1 23.0±3.8 13.1±3.1 17.2±3.4 19.7±3.6 20.5±3.7 20.5 -11.2
SmolVLM2-2.2B-Instruct 38.7±4.4 23.8±3.9 25.4±3.9 29.5±4.1 34.4±4.3 30.3±4.2 30.4 -10.0
@@
In Table˜4, we compare PISA-Bench with commonly known benchmarks MMLU (hendrycks2021measuringmassivemultitasklanguage) and MMMU (yue2024mmmu). We observe that our benchmark is significantly more difficult compared to MMLU for the models investigated, e.g., we find a performance of 68.5 on MMLU using Qwen2.5-VL (7B), whereas the corresponding performance on PISA-Bench is only 52.4. However, we note that MMLU is a text-only benchmark, emphasizing that modalities other than text is more difficult for current language models.
@@ 6.  Related Work
Recent progress in large vision-language models (LVLMs) has been closely tied to the development of evaluation benchmarks. Early benchmarks primarily assessed visual perception and image understanding fu2023mme; MMBench; liu2023visualinstructiontuning; meng2024mmiumultimodalmultiimageunderstanding, often restricting evaluation to multiple-choice or short-form VQA tasks. More recently, these benchmarks have been extended to more general areas such as cognition and reasoning, for example, benchmarks such as M3Exam (zhang2023m3exammultilingualmultimodalmultilevel), SOK-Bench (wang2024sokbenchsituatedvideoreasoning), MathVista (lu2024mathvistaevaluatingmathematicalreasoning), or VL-ICL-Bench (zong2025vliclbenchdevildetails). Further, benchmarks like MMDU liu2024mmdu, adopt open-ended questions combined with LLM-as-a-judge evaluation, while MMMU-pro yue2024mmmu propose unifying text and images into a single visual representation.
@@
In the domain of multilingual benchmarks, authors often begin with English datasets and translate them into other languages, such as XNLI (conneau-etal-2018-xnli) or XCOPA (ponti-etal-2020-xcopa). More recent works utilize multilingual language models to translate the original test set, such as HumanEval-XL (peng2024humanevalxlmultilingualcodegeneration) or mHumanEval (raihan-etal-2025-mhumaneval). However, while this approach provides broad coverage, it may propagate cultural or linguistic biases through the translation using language models shi2022language. P-MMEval (zhang2024p) and BenchMAX (huang2025benchmax) address this issue by using parallel corpora to fairly assess cross-lingual capabilities, disentangling cultural knowledge from a language model’s translation ability.
@@
Many multilingual LVLM benchmarks evaluate the general natural language and image understanding capabilities of vision-language models such as xGQA (pfeiffer2021xgqa), GEM (su2021gemgeneralevaluationbenchmark), and MaXM (changpinyo2022maxm). More recent benchmarks extend this line of work to reasoning and cognition, including M3Exam (zhang2023m3exammultilingualmultimodalmultilevel), EXAM-V (das2024exams), or M5-VGR (schneider2024m5). Others focus on culture-specific reasoning, such as ALM-bench (vayani2024all) and CVQA (romero2024cvqa), or domain-specific reasoning, such as medical reasoning in WorldMedQA-V(matos2024worldmedqavmultilingualmultimodalmedical). To mitigate the challenges of translating datasets into target languages, parallel benchmarks such as M4U (wang2024m4u), PM4Bench (gao2025pm4benchparallelmultilingualmultimodal), and XT-VQA (yu2024cross) have been proposed.
@@ 7.  Conclusion
Acknowledgments
This paper has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101189745 (HIVEMIND) and the German Research Foundation (DFG) project NFDI4DS under Grant No.: 460234259. We would also like to thank Yuan Li and her team from the Institute of German Studies at Zhejiang University, as well as Martin Courtois, Carlos Franzreb, and Alessandro Noveta, for annotating the dataset and verifying the translations.
Alan Akbik and Patrick Haller are supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Emmy Noether grant “Eidetic Representations of Natural Language” (project number 448414230). Further, Alan Akbik is supported under Germany’s Excellence Strategy “Science of Intelligence” (EXC 2002/1, project number 390523135). Jonas Golde is supported by the Bundesministerium für Bildung und Forschung (BMBF) as part of the project “FewTuRe” (project number 01IS24020).